# How turbulence increases the bubble-particle collision rate

### v1

@@ Abstract
Imposed bubble velocity.—We start from the discussion of the bubble with an imposed velocity. In Fig. 1(b) we demonstrate that the impact of turbulent fluctuations on the mean streamlines is insignificant for the present parameters. In particular in the incoming flow, which determines the collision rate, the differences are very small, lending support to our modelling approach. More noticeable differences in the wake region are a consequence of turbulence disrupting the symmetric recirculation pattern behind the bubble. However, the influence of turbulence on the collision process is distinct, as becomes evident from the movie [36] and Fig. 1(c), where trajectories of colliding particles are shown. These trajectories originate from a cone-shaped region that is notably larger than the corresponding collision radius $r_{c}$, based on the grazing trajectory in quiescent flow. Correspondingly, also the collision angle relative to the direction of the imposed bubble velocity is found to be wider in the turbulent flow [36]. Nevertheless, the collision trajectories appear to follow straight paths towards the bubble, which is consistent with our modelling assumptions.
@@
For quiescent flow, the dependence of $E_{c}$ normalised by the result for tracer particles $E_{c}(Re_{b},0)$, on $Re_{b}$ and $St_{p}$ is presented in Fig. 2(a). For small $St_{p}$, the inertial effect is negligible and the particles follow the flow streamlines, so that interception is the dominant factor determining the number of collisions in this range [47]. Analytical predictions based on flow streamlines in quiescent flow indicate that $\Gamma$ scales with $(r_{p}/r_{b})^{2}$[14, 48] in this case. However, as particle inertia becomes more pronounced, particles deviate from the flow streamlines and collide on the bubble even if their initial position is outside the grazing trajectory of the inertialess particle. This leads to a higher collision rate and explains the rapid increase in $E_{c}$ as $St_{p}$ approaches 0.1, beyond which the inertial effect dominates. At very large $St_{p}$, the particles are barely influenced by the flow such that $E_{c}$ approaches $(1+r_{p}/r_{b})^{2}$. To be able to evaluate $E_{c}(Re_{b},St_{p})$ analytically, we employ an empirical expression, which is the sum of its two contributions: interceptional collision $E_{i}$ and the collision associated with particle inertia[14]:
@@
Figure 3: (a) Contour lines of $P(r,l)$ from simulations (solid) and model (dashed lines) for $r_{p}/r_{b}=0.05$ and $St_{p}=0.04$. The grey shaded area indicates the region inside the grazing trajectory in quiescent flow. Normalized collision kernel for the freely rising bubble at (b) $Re_{\lambda}=32$, $Re_{b}=110$ and (c) $Re_{\lambda}=64$, $Re_{b}=150$. The inset in (b) compares the auto-correlation function of $U_{b}^{\prime}$ for the imposed velocity ($Re_{\lambda}=32,~{}Re_{b}=120$) and freely rising bubbles ($Re_{\lambda}=32,~{}Re_{b}=110$), respectively. The inset in (c) shows the normalized mean incoming particle density as a function of $St_{p}$ at $Re_{\lambda}=64$. (d) Sketch in the bubble reference frame, showing how fluctuations $U_{b}^{\prime}$ during the interaction effectively enlarge the particle collision radius.
As shown for the case with constant bubble velocity in Fig. 2(b), the general trends of $\Gamma$ in turbulence, in particular the increase with increasing $St_{p}$ for all particle sizes are consistent with those observed in quiescent flow. In this simplified configuration, the PDF of the $U_{b}^{\prime}$ required to evaluate our model (Eq. 1), can be obtained by combining the constant bubble rise velocity with the Gaussian distribution of turbulent velocity fluctuations [36]. The collision kernel predicted in this way is in excellent agreement with the simulations. This is further confirmed in Fig. 2(c), where we scrutinize the result by plotting it relative to the collision kernel at the same bubble velocity in quiescent flow as $\Gamma^{*}=\Gamma/\Gamma^{q}$. In this way, it also becomes clear that turbulent flow significantly enhances the collision kernel. Two interesting aspects should be underscored: Firstly, $\Gamma^{*}$ surpasses 1 when inertia is negligible ($St_{p}\ll 1$), indicating that turbulent fluctuations enhance interceptional bubble-particle collision. Secondly, the collision enhancement is not uniform across the considered range of $St_{p}$, suggesting that the combined influence of turbulence and particle inertia leads to further amplification of the collision rate. The collision enhancement can reach approximately 100% for particles with a size ratio of $r_{p}/r_{b}=0.025$. For larger particles, the maximum collision enhancement is lower, though the peak still occurs at a similar value of $St_{p}\approx 0.1$.
@@
In addition, there is an inertial effect as a change in $U_{b}^{\prime}$ also changes $St_{p}^{\prime}$. The strongly nonlinear dependence of $E_{c}$ on $St_{p}^{\prime}$, especially in the intermediate range $St_{p}\approx 0.1$, leads to an asymmetry response to positive and negative velocity fluctuations. This is demonstrated by the scaled PDF of $St_{p}^{\prime}$, $F(St_{p}^{\prime})=E_{c}(St_{p}^{\prime})PDF(St_{p}^{\prime})$, shown in the inset of Fig. 2(a). For low and high $St_{p}$, the dependence of $F(St_{p}^{\prime})$ on $St_{p}^{\prime}$ is almost symmetric. However, for an intermediate value of $St_{p}\simeq 0.1$, the contribution to collisions from fluctuating $St_{p}^{\prime}$ exhibits a positive bias, which explains the strongly enhanced turbulent collision rate in this range and the non-monotonic dependence on $St_{p}$. Due to the higher interceptional collision efficiency, the increase in the inertial range and hence the turbulent enhancement is less pronounced for larger size ratios $r_{p}/r_{b}$.
Another way to validate the model is to consider the spatial distribution of the collision probability $P(r,~{}l)=N_{c}(r,l)/N(r,l)$, where $r$ and $l$ are distances perpendicular and along the bubble velocity direction, respectively, (see Fig. 1(c)) and $N_{c}(r,~{}l)$ measures the number of colliding particles out of the total number of particles $N(r,l)$ passing through position $(r,l)$ in a given time interval. $P(r,~{}l)$ can be predicted from the model by superimposing the appropriate binary distributions for quiescent flow [36]. The result of doing so is again in excellent agreement with the data as shown in Fig. 3(a), confirming that the basic physical processes are well represented in the model. Consistent with Fig. 1(c), the impact of turbulence is clearly evident leading to a much wider distribution of $P(r,~{}l)$ compared to the binary distribution in quiescent flow (grey shading).
@@
Freely rising bubble.— Having established the general suitability of the model to capture the relevant turbulence effects on the collision rate, we now turn to the more realistic case of a freely rising bubble. The corresponding results in terms of $\Gamma$ are shown in Fig. 3(b,c) for different $Re_{\lambda}$. In these cases, the bubble-slip velocity pdf $f(U_{b}^{\prime})$ required as input for the model is obtained by measuring the fluid velocity around the bubble [36, 49].
@@
The general agreement between $\Gamma$ predicted by the model (dashed lines in Fig. 3(b, c) and the data remains very good also for the free rising cases. Small difference arise at low $St_{p}$, where the model is found to slightly underpredict the simulation result. This can be explained by the correlation time of $U_{b}^{\prime}$, which is shorter for free rising bubbles compared to the imposed velocity case (see inset in Fig. 3(b)). The resulting changes in $U_{b}^{\prime}$ during the bubble-particle interaction cause the particle trajectory to fluctuate in the bubble frame of reference (see Fig. 3(d)). This increases the effective collision radius of the particle to $r_{p}+d$, where $d$ is a measure of the drift from the original particle trajectory. The relevant time and velocity scales for this drift are $\tau_{f}$ and $u_{\eta}$, respectively, such that $d\sim\tau_{f}u_{\eta}$ in analogy to Taylor dispersion in the ballistic regime [50]. Indeed, we find that using $d=0.06\tau_{f}u_{\eta}$ results in good agreement with our data across different $Re_{\lambda}$ and for different $r_{p}$, as shown by the dash-dotted lines in Fig. 3(b,c). This effect is only relevant at $St_{p}\lessapprox 0.1$ and becomes negligible once $r_{c}\gg d$ due to the inertial effect at larger $St_{p}$.
@@
Another finding is that the incoming particle number density can differ from the global value. This is caused by clustering of bubbles and particles in different regions of the flow leading to segregation [8]. As a result $\langle n_{p}^{\prime}\rangle/n_{p}$ shown in the inset of Fig.  3(c) decreases as $St_{p}$ increases reaching a minimum value of about 0.8 for $St_{p}\approx 1$, where clustering effects are known to be strongest. The segregation effect explains why the inertial limit $(1+r_{p}/r_{b})^{2}$ at high $St_{p}$ is not reached in the simulations and in the model, where this effect is accounted for.
@@
Conclusions.—We have elucidated the relevant mechanisms governing the collision rate of inertial particles with a finite-size bubble in turbulence. We demonstrated that for the investigated practical conditions, inertial effects induced by the flow around the bubble are the dominant effect. The non-linear dependence of these effects on the bubble slip velocity leads to an increase of the collision rate in turbulence of up to 100% at $St_{p}\approx 0.1$ compared to quiescent flow. An additional increase in the turbulent collision rate is due to the short temporal correlation of the bubble slip velocity in the free-rising case. The effect of the resulting fluctuations during the bubble-particle interaction can be captured by an increase in the effective collision radius of the particle and is mostly relevant in the tracer limit for $St_{p}\lessapprox 0.1$. Segregation of bubbles and particles in turbulence reduces the particle density encountered by the bubble and and hence the collision rate by up to 20% at $St_{p}\approx 1$. Remarkably, the effect of turbulence induced motion of the particles was found to have negligible impact on the collision process throughout. The developed frozen turbulence model provides a physically consistent framework that can be easily extended to a full collision model (by combining it with a prediction for $f(U_{b}^{\prime})$). The approach is also transferable to other conditions, such as more complex shapes (by adopting a different parameterisation of $E_{c}$) and thus offers a more general relevance for collisions with finite-size objects in turbulence.

### v2

@@ IV Results
We start from the discussion of the bubble with an imposed velocity. In Fig. 1(c) we demonstrate that the impact of turbulent fluctuations on the mean streamlines is insignificant for the present parameters. In particular in the incoming flow, which determines the collision rate, the differences are very small, lending support to our modelling approach. More noticeable differences in the wake region are a consequence of turbulence disrupting the symmetric recirculation pattern behind the bubble. However, the influence of turbulence on the collision process is distinct, as becomes evident from the supplementary movie and Fig. 1(d), where trajectories of colliding particles are shown. These trajectories originate from a cone-shaped region that is notably larger than the corresponding collision radius $r_{c}$, based on the grazing trajectory in quiescent flow. Correspondingly, also the collision angle relative to the direction of the imposed bubble velocity is found to be wider in the turbulent flow (see Appendix A). Nevertheless, the collision trajectories appear to follow straight paths towards the bubble, and the colliding particle stream exhibits a band-like pattern, as shown in Fig. 1(e), both of which are consistent with our modeling assumptions.
Figure 2: (a) Normalised $E_{c}$ as a function of $St_{p}$ in quiescent flow for various $Re_{b}$ compared to the fits according to Eq. (9) (dashed lines). (b) Dimensionless collision kernel vs. $St_{p}$ at $\overline{Re_{b}}=120$ for bubble with imposed velocity in HIT (symbols) and model (dashed lines). (c) Turbulent collision kernel relative to that in quiescent flow for the imposed velocity bubble. Error bars represent fluctuations between subsets of the data. (d) Scaled PDF of $St_{p}^{\prime}$ as a function of $St_{p}^{\prime}$ for different $St_{p}$.
@@
For quiescent flow, the dependence of $E_{c}$ normalised by the result for tracer particles $E_{c}(Re_{b},0)$, on $Re_{b}$ and $St_{p}$ is presented in Fig. 2(a). For very small $St_{p}$, the inertial effect is negligible and the particles follow the flow streamlines, so that interception is the dominant factor determining the number of collisions in this range [26]. Analytical predictions based on flow streamlines in quiescent flow indicate that $\Gamma$ scales with $(r_{p}/r_{b})^{2}$ [38] in this case. However, as particle inertia becomes more pronounced, particles deviate from the flow streamlines and collide on the bubble even if their initial position is outside the grazing trajectory of the inertialess particle. This leads to a higher collision rate and explains the rapid increase in $E_{c}$ as $St_{p}$ approaches 0.1, beyond which the inertial effect dominates. At very large $St_{p}$, the particles are barely influenced by the flow such that $E_{c}$ approaches $(1+r_{p}/r_{b})^{2}$. To be able to evaluate $E_{c}(Re_{b},St_{p})$ analytically, we employ an empirical expression, which is the sum of its two contributions: interceptional collision $E_{i}$ and the collision associated with particle inertia[8]:
@@
As shown for the case with constant bubble velocity in Fig. 2(b), the general trends of $\Gamma$ in turbulence, in particular the increase with increasing $St_{p}$ for all particle sizes are consistent with those observed in quiescent flow. In this simplified configuration, the PDF of the $U_{b}^{\prime}$ required to evaluate our model (Eq. 2), can be obtained by combining the constant bubble rise velocity with the Gaussian distribution of turbulent velocity fluctuations. The collision kernel predicted in this way is in excellent agreement with the simulations. This is further confirmed in Fig. 2(c), where we scrutinize the result by plotting it relative to the collision kernel at the same bubble velocity in quiescent flow as $\Gamma^{*}=\Gamma/\Gamma^{q}$. In this way, it also becomes clear that turbulent flow significantly enhances the collision kernel. Here, two interesting aspects should be underscored: Firstly, $\Gamma^{*}$ surpasses 1 when inertia is negligible ($St_{p}\ll 1$), indicating that turbulent fluctuations enhance interceptional bubble-particle collision. Secondly, the collision enhancement is not uniform across the considered range of $St_{p}$, suggesting that the combined influence of turbulence and particle inertia leads to further amplification of the collision rate. The collision enhancement can reach approximately 100% for particles with a size ratio of $r_{p}/r_{b}=0.025$. For larger particles, the maximum collision enhancement is lower, though the peak still occurs at a similar value of $St_{p}\approx 0.1$.
@@
In addition, there is an inertial effect as a change in $U_{b}^{\prime}$ also changes $St_{p}^{\prime}$. The strongly nonlinear dependence of $E_{c}$ on $St_{p}^{\prime}$, especially in the intermediate range $St_{p}\approx 0.1$, leads to an asymmetric response to positive and negative velocity fluctuations. This is demonstrated by the scaled PDF of $St_{p}^{\prime}$, $F(St_{p}^{\prime})=E_{c}(St_{p}^{\prime})PDF(St_{p}^{\prime})$, shown in Figure 2(d). For low and high $St_{p}$, the dependence of $F(St_{p}^{\prime})$ on $St_{p}^{\prime}$ is almost symmetric. However, for an intermediate value of $St_{p}\simeq 0.1$, the contribution to collisions from fluctuating $St_{p}^{\prime}$ exhibits a positive bias, which explains the strongly enhanced turbulent collision rate in this range and the non-monotonic dependence on $St_{p}$. Due to the higher interceptional collision efficiency, the increase in the inertial range and hence the turbulent enhancement is less pronounced for larger size ratios $r_{p}/r_{b}$.
Figure 3: (a) Sketch of the bubble-particle collision with temporary slip velocity $\mathbf{U}_{b}^{\prime}$. The blue shaded region indicates the binary function $S(r,l,\mathbf{U}_{b}^{\prime})$, which is the projection of the collision tube on the $r-l$ plane. (b) Sketch of the cross section (grey region) between the collision tube and the plane $l$ in the view along $l$ axis. $\theta(r,l,\mathbf{U}_{b}^{\prime})$ indicates the radian of the arc that occupied by the grey region in the circle of radius $r$, which is used measure the term $G(r,l,\mathbf{U}_{b}^{\prime})$. (c) and (d): Contour lines of $P(r,l)$ from simulations (solid) and model (dashed lines) for $(r_{p}/r_{b}=0.05,St_{p}=0.04)$ and $(r_{p}/r_{b}=0.1,St_{p}=0.37)$, respectively. (d) Sketch of the collision probability under different bubble slip velocities. The region with more overlaps corresponds to the one with higher collision probability.
Another way to validate the model is to consider the spatial distribution of the collision probability $P(r,~{}l)$, where $r$ and $l$ are distances perpendicular and along the bubble velocity direction, respectively, (see Fig. 3(a)). In quiescent flow, the collision process is deterministic and the collisions probability is a binary function, which has a tube shape based on the grazing trajectory (see the shaded region in Fig. 3(a)). We represent this using the binary function $S(r,l,;\mathbf{U}_{b})$, which is equal to 1 inside the collision tube and zero otherwise, as shown in Figure 3(a). Practically, $S$ can be determined from the grazing trajectory. Based on the ‘frozen turbulence’ assumption, the collision probability in turbulence can be predicted as the superposition of the binary probability distributions corresponding to the instantaneous slip velocity $\mathbf{U}_{b}^{\prime}$. This leads to
$P(r,l)=\int f(\mathbf{U}_{b}^{\prime})S(r,l;\mathbf{U}_{b}^{\prime})G(r,l,\mathbf{U}_{b}^{\prime}){\rm d}\mathbf{U}_{b}^{\prime}.$ (10)
Note that the additional factor $G(r,l,\mathbf{U}_{b}^{\prime})=\theta(r,l,\mathbf{U}_{b}^{\prime})/2\pi$ in Eq. (10) is a geometrical coefficient related to the azimuthal integration required for the projection onto the two-dimensional $(r,l)$ space. It measures the fraction of the circle with radius $r$ that falls inside the collision tube at distance $l$ as illustrated in figure 3(b). The result of model prediction according to Eq. (10) is again in excellent agreement with the data as shown in Fig. 3(c) and (d), confirming that the basic physical processes are well represented in the model. Consistent with Fig. 1(d), the impact of turbulence is clearly evident leading to a much wider distribution of $P(r,~{}l)$ compared to the binary distribution in quiescent flow. This effect is especially pronounced if $St_{p}$ is low as is the case for Fig. 3 (c). The collision efficiency is low and the associated collision stream tube slender for this case, such that fluctuations in the instantaneous bubble slip direction lead to low values of $P(r,~{}l)$ even close to the bubble surface, as is illustrated in Fig. 3 (e). Since the collision efficiency increases for larger $St_{p}$ and the collision tube widens, this effect becomes less strong and the values of the collision probability close to the bubble are much higher in the plot for $St_{p}=0.37$ in Fig. 3 (d).
Figure 4: Normalized collision kernel for the freely rising bubble at (a) $Re_{\lambda}=32$, $\overline{Re_{b}}=110$ and (b) $Re_{\lambda}=64$, $\overline{Re_{b}}=150$. (c) Comparison of auto-correlation function of $U_{b}^{\prime}$ for the imposed velocity ($Re_{\lambda}=32,~{}\overline{Re_{b}}=120$) and freely rising bubbles ($Re_{\lambda}=32,~{}\overline{Re_{b}}=110$), respectively. (d)) shows the normalized mean incoming particle number density as a function of $St_{p}$ at $Re_{\lambda}=64$. (e) Sketch in the bubble reference frame, showing how fluctuations $U_{b}^{\prime}$ during the interaction effectively enlarge the particle collision radius.
@@
Having established the general suitability of the model to capture the relevant turbulence effects on the collision rate, we now turn to the more realistic case of a freely rising bubble. The corresponding results in terms of $\Gamma$ are shown in Fig. 4(a,b) for different $Re_{\lambda}$. In these cases, the bubble-slip velocity pdf $f(U_{b}^{\prime})$ required as input for the model is measured by averaging the fluid velocity located on a spherical surface of radius $3r_{b}$ that is centred at the bubble’s center position [39]. The size of the spherical surface is chosen in a way that the fluid velocity is not significantly influenced by the presence of the bubble boundary. The radius $3r_{b}$ of the spherical surface is tested in the case of uniform flow past a fixed sphere [39], which results in a measured fluid velocity approximately $90\%$ of the incoming flow velocity.
@@
The general agreement between $\Gamma$ predicted by the model (dashed lines in Fig. 4(a, b) and the data remains good also for the free rising cases. Difference arises at low $St_{p}$, where the model is found to underpredict the simulation result. This can be explained by the correlation time of $U_{b}^{\prime}$, which is shorter for free rising bubbles compared to the imposed velocity case (see Fig. 4(c)). The resulting changes in $U_{b}^{\prime}$ during the bubble-particle interaction cause the particle trajectory to fluctuate in the bubble frame of reference (see the sketch in Fig. 4(e)). This increases the effective collision radius of the particle to $r_{p}+d$, where $d$ is a measure of the drift from the original particle trajectory. The relevant time and velocity scales for this drift are $\tau_{f}$ and $u_{\eta}$, respectively, such that $d\sim\tau_{f}u_{\eta}$ in analogy to Taylor dispersion in the ballistic regime [40]. Indeed, we find that using $d=0.06\tau_{f}u_{\eta}$ results in good agreement with our data across different $Re_{\lambda}$ and for different $r_{p}$, as shown by the dash-dotted lines in Fig. 4(a,b). This effect is only relevant at $St_{p}\lessapprox 0.1$ and becomes negligible once $r_{c}\gg d$ due to the inertial effect at larger $St_{p}$.
@@
Another finding is that the incoming particle number density can differ from the global value. This is caused by clustering of bubbles and particles in different regions of the flow leading to segregation [7]. As a result $\langle n_{p}^{\prime}\rangle/n_{p}$ shown in Fig.  4(d) decreases as $St_{p}$ increases reaching a minimum value of about 0.8 for $St_{p}\approx 1$, where clustering effects are known to be strongest. The segregation effect explains why the inertial limit $(1+r_{p}/r_{b})^{2}$ at high $St_{p}$ is not reached in the simulations and in the model, where this effect is accounted for by multiplying with the factor $\langle n_{p}^{\prime}\rangle/n_{p}$ obtained from the simulations.
V Conclusions
@@
We have elucidated the relevant mechanisms governing the collision rate of inertial particles with a finite-size bubble in turbulence. We demonstrated that for the investigated practical conditions, inertial effects induced by the flow around the bubble are the dominant effect. The non-linear dependence of these effects on the bubble slip velocity leads to an increase of the collision rate in turbulence of up to 100% at $St_{p}\approx 0.1$ compared to quiescent flow. An additional increase in the turbulent collision rate is due to the short temporal correlation of the bubble slip velocity in the free-rising case. The effect of the resulting fluctuations during the bubble-particle interaction can be captured by an increase in the effective collision radius of the particle and is mostly relevant in the tracer limit for $St_{p}\lessapprox 0.1$. Segregation of bubbles and particles in turbulence reduces the particle density encountered by the bubble and and hence the collision rate by up to 20% at $St_{p}\approx 1$. Remarkably, the effect of turbulence induced motion of the particles was found to have negligible impact during the transient bubble-particle interaction, whereas the incoming particle number density is reduced for particles with large Stokes number. The developed frozen turbulence model provides a physically consistent framework that can be easily extended to a full collision model (by combining it with a prediction for $f(U_{b}^{\prime})$). The approach is also transferable to other conditions, such as more complex shapes (by adopting a different parameterisation of $E_{c}$) and thus offers a more general relevance for collisions with finite-size objects in turbulence.
@@ Acknowledgements.
Declaration of Interests. The authors report no conflict of interest.
Appendix A
Figure 5: (a) Mean collision angle $\langle\theta\rangle$ as a function of $St_{p}$. Sketches in panel (b) and panel (c) illustrate the mechanism that $\langle\theta\rangle$ declines/increase as increasing inertia when $St_{p}$ is low/large in quiescent flow. The solid lines denote the streamlines which the particles originally stay and the dashed lines are the particle trajectories. Panel (d-f) show the PDF of collision angle $f(\theta)$ scaled by the normalized collision kernel for particle size $r_{p}/r_{b}=0.025,~{}0.05,~{}0.1$ in HIT (solid lines) and quiescent flow (dashed lines).
As we observe that the trajectories of collided particles in HIT are significantly different from those in quiescent flow, it is interesting to study how turbulence affects the collision angle. In Fig. 5 we investigate the collision angle $\theta$ for the case where the bubble velocity is imposed. Here, $\theta$ denotes the angle between the direction of the imposed bubble velocity and the vector of collision position in the bubble frame. We observe that the mean collision angle $\langle\theta\rangle$ for the turbulence flow is higher than that for the case of quiescent flow. However, $\langle\theta\rangle$ shows a similar dependence on $St_{p}$ for both cases, where $\langle\theta\rangle$ first decreases as $St_{p}$ and rises after. This can be explained by inertial effects on the particles as illustrated in Fig. 5 (b) and (c) for the deterministic case. At low $St_{p}$, $r_{c}$ is small. Consequently, particles collide on the bubble front region, where the streamlines are strongly curved. Therefore particles trajectories significantly deviate from the streamlines as $St_{p}$ increases. As a consequence, a higher inertia particles collide at a smaller collision angle, as illustrated in Fig. 5 (b). When $St_{p}$ surpasses the critical value, $r_{c}$ becomes larger. In this case, increasing particle inertia leads to a higher collision rate as well as to a higher collision angle, which is illustrated in Fig. 5 (c). These general trends for $\langle\theta\rangle$ are retained in the turbulent case, which is in line with our approach of representing the instantaneous collision process by the deterministic case. Fig. 5 (d-f) show the scaled PDF of collision angle $\theta$ for different particle size, respectively. The scaled PDF illustrates the distribution of the collision angle, as well as where the collision difference occur between HIT and quiescent flow. The distributions of collision angle indicate that the collisions are more likely to take place at intermediate angle. Moreover, we observe that the distribution extend to a wider range in HIT and more collisions occurs, which are consistent with the observations for $\langle\theta\rangle$.